# CoMoE: Contrastive Representation for Mixture-of-Experts in Parameter-Efficient Fine-tuning

### v1

@@ 2 $s_{\mathrm{pos}}\leftarrow(q\cdot P^{\top})/\tau$
    Positive set. Aggregate the remaining $(k-1)$ expert representations, excluding the one indexed by $T[r]$ from the index set $T$, into a set $P$, applying normalization:
@@ Appendix H Model efficiency and Computational budget
To evaluate the model efficiency and computational budget of CoMoE, we compare it with the baselines(LoRA, DoRA, MixLoRA, OMOE) in three aspects: inference latency, memory cost, and training time. We base our evaluation on the following three metrics: (a) the inference time required for generating responses (ms), (b) the GPU memory cost (MiB), and (c) the training time in multi-task settings (h). In the multi-task setting, ARC-e, ARC-c, BoolQ, and OBQA are trained simultaneously. The results are provided in Table 8. From Table 8, we observe that compared to the well-performing MixLoRA, CoMoE achieves a 10% improvement in inference efficiency while reducing GPU memory usage by 465 MiB. In terms of training time, CoMoE requires 3.5 hours of training on an A6000 GPU under the multi-task setting. Although CoMoE increases the training burden, it does not compromise inference efficiency and simultaneously enhances model performance.
@@ Appendix I Additional Visualization of Representations under Different Datasets
CoMoE-LoRA 1.45% 79.0 55.5 72.9 80.0 83.6
CoMoE-DoRA 1.45% 78.7 57.0 72.9 78.4 83.8

### v2

@@ 2 $s_{\mathrm{pos}}\leftarrow(q\cdot P^{\top})/\tau$
    Positive set. Aggregate the remaining (k1)(k-1) expert representations, excluding the one indexed by $T[r]$ from the index set $T$, into a set $P$, applying normalization:
@@ Appendix H Model efficiency and Computational budget
To evaluate the model efficiency and computational budget of CoMoE, we compare it with the baselines(LoRA, DoRA, MixLoRA, OMOE) in three aspects: inference latency, memory cost, and training time. We base our evaluation on the following three metrics: (a) the inference time required for generating responses (ms), (b) the GPU memory cost (MiB), and (c) the training time in multi-task settings (h). In the multi-task setting, ARC-e, ARC-c, BoolQ, and OBQA are trained simultaneously. The results are provided in Table 9. From Table 9, we observe that compared to the well-performing MixLoRA, CoMoE achieves a 10% improvement in inference efficiency while reducing GPU memory usage by 465 MiB. In terms of training time, CoMoE requires 3.5 hours of training on an A6000 GPU under the multi-task setting. Although CoMoE increases the training burden, it does not compromise inference efficiency and simultaneously enhances model performance. Furthermore, we conducted experiments to investigate the relationship between training time and the number of experts, as shown in Table 10. In our proposed method, negative samples are drawn from the set of inactive experts. As a result, the training cost increases linearly with the number of experts, with a time complexity of $O(n)$. To mitigate this computational overhead, we introduce a fixed-size sampling strategy, wherein a constant number of negative samples are randomly selected from the inactive experts at each training step. This reduces the complexity from $O(n)$ to constant time $O(1)$, rendering it independent of the total number of experts. The results of this strategy are summarized in Table 11. Theoretically, this approach introduces a looser lower bound (see Theorem 1); however, empirical results show no degradation in performance, as evidenced by the comparable average accuracy between Table 10 and Table 11.
@@
Number of Experts 4 5 6 7 8
Training Time (h) 7.5 9.0 10.8 12.0 13.0
Average Accuracy 0.7248 0.7309 0.7248 0.7381 0.7239
Table 10: Training time and accuracy with varying number of experts. Training time grows linearly with expert count.
Number of Experts 4 5 6 7 8
Training Time (h) 7.5 7.5 7.5 7.6 7.6
Average Accuracy 0.7424 0.7357 0.7311 0.7305 0.7319
Table 11: Training time and accuracy after applying fixed-size negative sampling. Training time remains constant regardless of expert count.
@@ Appendix I Additional Visualization of Representations under Different Datasets
CoMoE-LoRA 1.45% 80.3 57.3 72.9 80.4 83.6
CoMoE-DoRA 1.45% 80.2 57.0 73.3 81.2 83.8
@@
Appendix J Analysis of Performance Differences Between Multi-task and Single-task Settings
CoMoE exhibits consistently stronger performance in multi-task setups compared to single-task training, as demonstrated in both math and reasoning benchmarks. This observation suggests that the expert specialization mechanism of CoMoE is particularly well-suited for heterogeneous data distributions. In contrast, under simple or narrow single-task settings (e.g., ARC-c or OBQA), models may overfit to the specific task distribution and potentially achieve better results through memorization rather than generalization. To further illustrate this difference, consider the following math problem:
“Eight more than four times the number of coffee mugs in the teacher’s lounge is 40. How many coffee mugs are in the teacher’s lounge?”
A model trained under the single-task setting responds:
“Assistant: Eight more than four times the number of coffee mugs in the teacher’s lounge is 40. That means four times the number of coffee mugs in the teacher’s lounge is 40 – 8 = 32. So, the number of coffee mugs in the teacher’s lounge is 32. The answer is 32.”
This output is incorrect—it skips formal equation modeling and instead relies on direct arithmetic subtraction, leading to an erroneous conclusion. This reflects a typical overfitting behavior, where the model learns shallow patterns from the training set rather than robust reasoning strategies. Such behavior is more likely in single-task regimes due to limited exposure to diverse reasoning formats. In contrast, the same problem posed to a multi-task trained model yields:
“Assistant: Eight more than four times the number of coffee mugs is 40. We can write this as 4x + 8 = 40. We can solve this equation to find x. 4x + 8 = 40. 4x = 40 – 8. 4x = 32. x = 32 / 4 = 8. The number of coffee mugs must be 8. The answer is 8.”
This response correctly models the algebraic structure and performs a step-by-step solution. It demonstrates stronger generalization and compositional reasoning ability, which we attribute to the diversity and inductive bias introduced by multi-task training. These findings underscore the advantage of CoMoE in multi-task settings, especially for problems requiring structured reasoning.
Appendix K An Intuitive Illustration of Expert Specialization
To clarify the intuition behind our expert specialization design, consider the following illustrative scenario: Alice, Bob, and Carol are working together to develop a robot. Initially, all three study mathematics, programming, and hardware simultaneously. However, their lack of focus leads to only moderate proficiency across all domains. Later, they adopt a collaborative specialization strategy: mathematics is primarily handled by Alice and Bob, programming by Bob and Carol, and hardware by Alice and Carol. In this arrangement, Alice takes the lead in mathematics with Bob assisting; Bob leads programming with Carol assisting; and Carol leads hardware assembly with Alice assisting.
This role assignment enables each person to develop deeper expertise in their respective areas. The result is a more effective and efficient collaboration, ultimately leading to the successful construction of the robot.
Translating this scenario into our model design, we employ contrastive learning to formalize the process of specialization. For example, we encourage Alice and Bob (experts) to focus on learning mathematical tasks while discouraging Carol’s involvement in that domain. This negative signal allows Carol to concentrate on mastering programming or hardware instead. In practice, such structured encouragement and discouragement drive the emergence of distinct and complementary expert roles within the model, enabling better generalization across diverse tasks.